OpenStudy (anonymous):

Find the indicated limit, if it exists. limit of f(x), f(x)={5x-8 x<0,} x->0 {|-4-x| x>0} Please explain D:

OpenStudy (anonymous):

@terenzreignz do you know how to do these D:

OpenStudy (anonymous):

\[find \lim_{x \rightarrow0-}f \left( x \right)=\lim_{x \rightarrow 0-}\left( 5x-8 \right)\]

OpenStudy (anonymous):

I don't understand? btw the 5x-8 and the abs value of -4-x are all in one parentheses

OpenStudy (anonymous):

\[\left| 4-x \right|=-\left( 4-x \right),if 4-x<0,or4<x or x>4\]\[\left| 4-x \right|=4-x if 4-x>0 or 4>x or x<4,\] \[we have tofind the limit as x \rightarrow 0 sotake f(x)=4-x\]